The height of a garage door measures four feet less than the door’s width. The area of the garage door is 165 square feet. Determine the height and width of the garage door.
i need the complete algebraic work for this or it won't be graded

Answer by mananth(16946)


The height of a garage door measures four feet less than the door’s width.
let door's width be w
height of door will be w-4
The area of the garage door is 165 square feet.
Area of door = L * W
165 = w(w-4)
165 =w^2-4w
w^2-4w -165 =0
w^2 -15w+11w-165=0
w(w-15) +11(w-15)=0
(w-15)(w+11)=0
w =15 0r -11
ignore negative, door cannot have negative width
width =15 ft.
length = 15-4 = 11 ft